# Large Language Models for Multilingual Vulnerability Detection: How Far Are We?

### v1

@@ Abstract.
Various deep learning-based approaches utilizing pre-trained language models (PLMs) have been proposed for automated vulnerability detection. With recent advancements in large language models (LLMs), several studies have begun exploring their application to vulnerability detection tasks. However, existing studies primarily focus on specific programming languages (e.g., C/C++) and function-level detection, leaving the strengths and weaknesses of PLMs and LLMs in multilingual and multi-granularity scenarios largely unexplored. To bridge this gap, we conduct a comprehensive fine-grained empirical study evaluating the effectiveness of state-of-the-art PLMs and LLMs for multilingual vulnerability detection. Using over 30,000 real-world vulnerability-fixing patches across seven programming languages, we systematically assess model performance at both the function-level and line-level. Our key findings indicate that GPT-4o, enhanced through instruction tuning and few-shot prompting, significantly outperforms all other evaluated models, including CodeT5P. Furthermore, the LLM-based approach demonstrates superior capability in detecting unique multilingual vulnerabilities, particularly excelling in identifying the most dangerous and high-severity vulnerabilities. These results underscore the promising potential of adopting LLMs for multilingual vulnerability detection at function-level and line-level, revealing their complementary strengths and substantial improvements over PLM approaches. This first empirical evaluation of PLMs and LLMs for multilingual vulnerability detection highlights LLMs’ value in addressing real-world software security challenges.
Multilingual Vulnerability, Vulnerability Detection, Large Language Model
journal: JACMarticle: 111copyright: acmcopyrightccs: Software and its engineering
1. Introduction
@@
Software vulnerabilities are flaws in code that pose significant risks to software systems, potentially leading to compromised sensitive information and system failures. To address these risks, researchers have developed various automated vulnerability detection (AVD) techniques. These techniques typically fall into three main categories: rule-based techniques, machine learning (ML)-based techniques, and deep learning (DL)-based techniques. Rule-based techniques (Gobbi and Kinder, 2023; Kim et al., 2019) and ML-based techniques (Grieco et al., 2016; Scandariato et al., 2014) require significant human expertise to design rules and features, yet often struggle to define effective patterns for identifying vulnerabilities. DL-based techniques often rely on Recurrent Neural Networks (RNNs) (Li et al., 2021b, a; Zou et al., 2021), Graph Neural Networks (GNNs) (Chakraborty et al., 2021; Zhou et al., 2019; Nguyen et al., 2022; Steenhoek et al., 2024), and Pre-trained Language Models (PLMs) (Hanif and Maffeis, 2022; Fu and Tantithamthavorn, 2022; Liu et al., 2024). These methods are relatively effective since they can learn vulnerability patterns in an end-to-end manner. Particularly, leveraging pre-trained knowledge, PLMs formulate vulnerability detection as a binary classification on code embedding, proven superior to RNNs and GNNs at detecting vulnerabilities (Steenhoek et al., 2023). Although PLMs show promise to some extent, these methods still face limitations in fully understanding code semantics for extracting representative or unseen vulnerability patterns (Ding et al., 2024).
@@
Large Language Models (LLMs) have recently garnered widespread attention and shown strong potential across a variety of code-related tasks (Hou et al., 2023). Several empirical studies (Fu et al., 2023a; Zhou et al., 2024a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c) have explored using LLMs for vulnerability detection. For instance, Fu et al. (2023a) found that GPT-3.5-Turbo and GPT-4 fail to detect single-language (i.e., C/C++) vulnerabilities at both function-level and line-level.  Yin (2024) revealed that ChatGPT, when using prompts alone, can be easily swayed to change vulnerability classifications, indicating low confidence. Furthermore,  Yin et al. (2024) demonstrated that while fine-tuned LLMs can detect vulnerabilities, they perform more weakly than PLMs. Moreover, some LLM-based AVD techniques have been proposed, including Vul-RAG (Du et al., 2024), MSIVD (Yang et al., 2024), and GRACE (Lu et al., 2024). These techniques demonstrate that LLMs are competitive with PLMs at detecting vulnerabilities in individual programming languages. For example, MSIVD applied multitask learning with LoRA to fine-tune the LLM and incorporated fine-tuned LLM embedding and Abstract Syntax Tree (AST) embedding. Vul-RAG checks code vulnerability by reasoning about vulnerability causes and fixing solutions from retrieved vulnerability knowledge, proving LLMs’ potential in detecting code vulnerabilities.
While LLMs-based and PLMs-based techniques have made significant progress on vulnerability detection, we identify three remaining limitations: (I) Focus on a single programming language. Most existing studies limit their vulnerability detection capabilities to C and C++, often relying on datasets such as CVEfixes (Bhandari et al., 2021) and Big-Vul (Fan et al., 2020). However, real-world development environments are inherently complex and often involve multiple programming languages simultaneously (Li et al., 2022), with existing studies further revealing the prevalence of vulnerabilities across diverse language ecosystems such as Python (Alfadel et al., 2023) and Go (Hu et al., 2024). To be practically applicable, research on AVD needs to expand beyond single-language vulnerability to address multilingual vulnerability. (II) Unclear comparative strengths of PLMs and LLMs on multilingual vulnerability scenario. PLMs and LLMs are considered two promising mainstream approaches. The relative advantages and limitations of PLMs versus LLMs in multilingual vulnerability detection remain underexplored. Additionally, existing LLM-based approaches often overlook the diversity of available LLMs and the variety of applicable prompting and fine-tuning strategies. This knowledge gap in understanding makes it difficult to select and improve models for specific tasks in multilingual contexts. (III) Lack of comprehensive evaluation across granularity levels. Vulnerability detection typically operates at two levels: function-level and line-level. Function-level detection determines whether a given function contains any vulnerabilities. However, the coarse granularity of this approach can hinder developers from efficiently identifying and addressing specific issues. In contrast, line-level detection aims to pinpoint the exact lines within a vulnerable function that are responsible for the vulnerability. Existing research (Zhou et al., 2024a) has primarily focused on function-level vulnerability detection in single-language settings, whereas the investigation of line-level multilingual vulnerability detection remains significantly more pressing. Additionally, multilingual contexts pose greater challenges for line-level vulnerability detection, as variations in syntax and semantics across programming languages increase the difficulty for language models to generalize effectively.
@@
Our key empirical findings are: (1) An appropriate LLM with instruction tuning and few-shot prompting demonstrates promising multilingual vulnerability detection performance. Specifically, for both function-level and line-level multilingual vulnerability detection, GPT-4o with instruction tuning and few-shot prompting demonstrates superior performance, achieving the highest accuracy (0.7196) at function-level compared to the best-performing CodeT5P (0.6037), and the best F1-score (0.6641) at line-level compared to other studied PLMs and LLMs. (2) Across different programming languages, the best-performing LLM achieves its highest function-level accuracy (0.8082) with Go and its lowest (0.6626) with Python. At the line-level, GPT-4o attains its highest F1-score (0.7815) on JavaScript, while its lowest F1-score (0.4348) is observed on C#. (3) Orthogonality analysis reveals that the best-performing LLM excels not only in uniquely correct detections but also shows higher accuracy in identifying the top-25 most dangerous CWE-IDs compared to PLMs. Base on CVSS severity, the top LLM also surpasses other studied PLMs and LLMs in detecting high-severity vulnerabilities (e.g., Critical and High level severity). (4) Furthermore, we investigate the impact of LLM size and compare effectiveness between reasoning and non-reasoning LLMs. The results show that the size of LLMs is not the decisive factor affecting performance on multilingual vulnerability detection, and the use of reasoning LLMs does not lead to significant improvements.
@@
  1. (1)
    This is the first empirical study to comprehensively investigate the effectiveness of PLMs and LLMs with various strategies on both function-level and line-level multilingual vulnerability detection.
  2. (2)
@@
  4. (4)
@@
    We have made the replication package publicly available on our homepage (rep, [n. d.]), including all used data, codes, and analysis details involved in our study.
@@
Paper Extension. This paper extends our prior work on a preliminary study (Yu et al., 2025), published as a research paper at the 1st International Workshop on Large Language Model Supply Chain Analysis, co-located with ISSTA’25. The key differences can be summarized as follows: (I) A broader range of studied models and prompting strategies. Specifically, we enhance PLMs by incorporating three OpenAI text-embedding models, and for LLMs, we explore two additional strategies: few-shot prompting and instruction tuning. (II) Incorporation of line-level detection. To provide a more comprehensive perspective, we extend the research scope beyond function-level detection to include line-level detection, enabling more detailed analysis of vulnerability locations. (III) Finer-grained analyses of detection performance. We conduct an in-depth analysis focusing on unique correct and incorrect detections, prediction tendencies for the Top-25 most dangerous CWE-IDs, and the effectiveness of detecting vulnerabilities across different severity levels. (IV) Exploration of effects of model size and architecture. We further investigate how the size of open-source LLMs and the use of reasoning-capable LLMs impact performance in multilingual vulnerability detection.
2. Related Work
2.1. Software Vulnerability
Software vulnerabilities are security flaws in software systems that malicious actors can exploit. These vulnerabilities can cause significant damage when left unpatched, particularly in widely-used software, leading to substantial financial losses (Bilge and Dumitras, [n. d.]). For example, the Log4Shell vulnerability (CVE-2021-44228) (Luttwak and Schindel, 2021) in the Log4j library, which allowed attackers to execute arbitrary Java code on servers and other systems, exposed approximately 93% of enterprise cloud environments to significant security risks. Two main classification systems in the community help categorize these issues: Common Weakness Enumeration (CWE) for identifying weakness types, and Common Vulnerability Exposure (CVE) for documenting specific vulnerability instances (CWE, 2024; CVE, 2024).
2.2. Automated Vulnerability Detection
Specifically,  Hanif and Maffeis (2022) introduced VulBERTa, a deep learning model that leverages a custom tokenization pipeline to pre-train RoBERTa (Liu et al., 2019) on real-world C/C++ code, enabling enhanced vulnerability detection through learned code syntax and semantics. LineVul (Fu and Tantithamthavorn, 2022) employs a BERT-based architecture (Devlin et al., 2019) to create code representations for detecting vulnerabilities at the function level and uses its attention mechanism to identify specific vulnerable lines. SVulD (Ni et al., 2023) trains a model to differentiate semantic representations of functions, irrespective of lexical similarity, for vulnerability detection, and provides natural language explanations to help developers intuitively understand the root causes of vulnerabilities.  Zhang et al. (2023b) decomposed the syntax-based Control Flow Graph (CFG) of a code snippet into multiple execution paths, extracted their representations through LLMs with intra- and inter-path attention mechanisms, and aggregated these representations to detect vulnerabilities.  Liu et al. (2024) pre-trains their PLM, PDBERT, by leveraging the Abstract Syntax Tree (AST) and Program Dependency Graph (PDG) of functions to predict statement-level control dependencies and token-level data dependencies. They then fine-tune the PLM for the vulnerability detection task. DeepDFA (Steenhoek et al., 2024) introduces a graph-based learning framework that incorporates graph embeddings with PLMs to enhance vulnerability detection performance. In this study, they use general PLMs as vulnerability detection baselines.
However, PLMs require vast amounts of data to uncover patterns of vulnerability, and their generalization ability may be limited when confronted with unseen data. Recently, LLMs that are pre-trained on extensive code corpora have demonstrated a remarkable ability to understand code semantics, showing significant performance improvements in code-related tasks (Hou et al., 2023). Several empirical studies (Fu et al., 2023a; Zhou et al., 2024a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c) have explored using LLMs for vulnerability detection.  Fu et al. (2023a) examined ChatGPT’s capability to detect vulnerabilities in C/C++ functions through a zero-shot prompt. They found that GPT-3.5-Turbo and GPT-4 fail to detect single-language vulnerabilities at both function-level and line-level.  Yin (2024) demonstrated that when relying solely on prompts, ChatGPT is highly susceptible to altering vulnerability classifications, reflecting low confidence in its assessments. Although the inclusion of contextual information enhances its accuracy, the model still struggles to reliably predict severity ratings for certain categories of CWEs. Vul-RAG (Du et al., 2024), a technique for vulnerability detection based on LLMs, constructs a comprehensive knowledge base of vulnerabilities, retrieves relevant information through semantic similarity, and utilizes reasoning to assess the presence of vulnerabilities in code. MSIVD (Yang et al., 2024) integrates a multitask sequence-to-sequence language model with program control flow graphs, which are encoded as graph neural networks, to enable sequence-to-classification vulnerability detection. GRACE (Lu et al., 2024) leverages graph structural information in the code and in-context learning to enhance the effectiveness of LLM-based software vulnerability detection.
3. Study Design
@@ 3.1. Dataset Preparation
Studied dataset. To evaluate PLM-based AVD approaches and LLMs for multilingual vulnerability detection, we utilize the REEF dataset (Wang et al., 2023b). This dataset contains 4,466 CVEs with 30,987 patches across seven programming languages, including comprehensive vulnerability information (CVE, CWE, CVSS, etc.) and project details (such as commit messages). REEF is constructed from real-world vulnerabilities sourced from the National Vulnerability Database (NVD) and Mend’s CVE list (WhiteSource, 2022), an open-source vulnerability database covering 2016-2023. At the function-level, the dataset consists of 6,957 functions written in C, 2,244 in C++, 1,529 in C#, 3,187 in Go, 6,207 in Java, 5,066 in JavaScript, and 5,797 in Python. Our study encompasses all seven programming languages supported by REEF: C, C++, C#, Go, Java, JavaScript, and Python.
@@
Data pre-processing. To adapt the REEF dataset for multilingual AVD tasks, our first step was to extract the vulnerable and non-vulnerable functions. We received commit data in raw format with patches. To analyze both the original and modified versions of vulnerable functions, we processed these patches using the Linux patch command. [^1] https://www.man7.org/linux/man-pages/man1/patch.1.html Following the commit data collection (Fan et al., 2020), we pre-processed the data by removing code comments to minimize bias. We then extracted function definition code using Tree-sitter (Brunsfeld, 2024), an efficient parser generator and incremental parsing library capable of analyzing multiple programming languages. To extract functions and their corresponding function-level and line-level labels, we iterated through function definitions in the pre-change file to match them with their post-change and pre-change function definitions. Specifically, the function names are used as identifiers to match and compare corresponding functions between the pre-change and post-change versions of the files. When multiple matching function definitions are found in the post-change file, we calculate the edit distance between the pre-change function and each post-change function. The pair with the smallest edit distance is then identified as the vulnerable and clean function pair. In other words, we define the pre-change function as the vulnerable function and the post-change function as the non-vulnerable function. Since many PLMs use the absolute position encoding, which is limited to the input length (e.g., 512), we filtered out functions that are greater than 512 in length. Finally, we obtained a total of 20,165 functions and corresponding labels. At the function-level, the dataset comprises 3,056 C, 1,792 C++, 427 C#, 2,905 Go, 3,235 Java, 5,468 JavaScript, and 3,282 Python functions. To obtain line-level labels, we first eliminated all non-vulnerable functions from the function-level dataset. We then applied DIFFLIB (Homepage, 2024) to match changed code lines by comparing each vulnerable function with its corresponding non-vulnerable version. DIFFLIB identifies three matching cases: (1) vulnerable functions fixed by adding code lines only, (2) vulnerable functions fixed by removing code lines only, and (3) vulnerable functions fixed by both removing and adding code lines. Since we need to locate specific line positions, only cases (2) and (3) can identify vulnerable line positions through deletion information. We use these two cases to obtain code line labels, i.e., line numbers and code lines, then filter out vulnerable functions where line labels cannot be extracted. Finally, we obtained a total of 8,134 vulnerable functions and the corresponding line labels. Finally, we obtained a total of 8,134 lines and corresponding labels. At the line-level, the dataset comprises 1,175 C, 651 C++, 151 C#, 919 Go, 1,375 Java, 2,496 JavaScript, and 1,367 Python functions.
Furthermore, to validate Tree-sitter’s robustness, we performed a sanity check. We randomly selected 377 samples, a sample ensuring a 95% confidence level with a 5% margin of error. The first author examined the function definitions between pre-change and post-change versions, verifying their consistency and alignment with the provided patch. Our manual inspection confirmed that Tree-sitter is reliable and generates no false positives.
@@
Table 1. Statistical summary of the function-level multilingual vulnerability dataset
@@
Table 2. Statistical summary of the line-level multilingual vulnerability dataset
@@
Construction of training and test datasets. Similar to Fu and Tantithamthavorn (2022), we used random sampling to divide the REEF dataset of function-level and line-level multilingual vulnerability into training, validation, and test sets at an 8:1:1 ratio. To maintain consistent proportions of data across languages, we first allocated the data of each language according to the predetermined ratio. We then combined these portions to create the final dataset. As shown in Table 1, the function-level multilingual vulnerability dataset contains 16,126, 2,013, and 2,026 functions and corresponding labels for training, validation, and testing over seven languages. As shown in Table 2, the line-level multilingual vulnerability dataset contains 6,513, 811, and 810 vulnerable functions and corresponding line labels for training, validation, and testing across the studied languages.
@@
Figure 1. The severity distribution of multilingual vulnerability dataset
3.2. Experimented Language Models

### v2

# Michael Fu

@@ Abstract
Various deep learning-based approaches utilizing pre-trained language models (PLMs) have been proposed for automated vulnerability detection. With recent advancements in large language models (LLMs), several studies have begun exploring their application to vulnerability detection tasks. However, existing studies primarily focus on specific programming languages (e.g., C/C++) and function-level detection, leaving the strengths and weaknesses of PLMs and LLMs in multilingual and multi-granularity scenarios largely unexplored. To bridge this gap, we conduct a comprehensive fine-grained empirical study evaluating the effectiveness of state-of-the-art PLMs and LLMs for multilingual vulnerability detection. Using over 30,000 real-world vulnerability-fixing patches across seven programming languages, we systematically assess model performance at both the function-level and line-level. Our key findings indicate that GPT-4o, enhanced through instruction tuning and few-shot prompting, significantly outperforms all other evaluated models, including CodeT5P. Furthermore, the LLM-based approach demonstrates superior capability in detecting unique multilingual vulnerabilities, particularly excelling in identifying the most dangerous and high-severity vulnerabilities. These results underscore the promising potential of adopting LLMs for multilingual vulnerability detection at function-level and line-level, revealing their complementary strengths and substantial improvements over PLM approaches. This empirical evaluation of PLMs and LLMs for multilingual vulnerability detection highlights LLMs’ value in addressing real-world software security challenges.
1 Introduction
@@
Software vulnerabilities are flaws in code that pose significant risks to software systems, potentially leading to compromised sensitive information and system failures. To address these risks, researchers have developed various automated vulnerability detection (AVD) techniques. These techniques typically fall into three main categories: rule-based techniques, machine learning (ML)-based techniques, and deep learning (DL)-based techniques. Rule-based techniques (Gobbi and Kinder, 2023; Kim et al., 2019) and ML-based techniques (Grieco et al., 2016; Scandariato et al., 2014) require significant human expertise to design rules and features, yet often struggle to define effective patterns for identifying vulnerabilities. DL-based techniques often rely on Recurrent Neural Networks (RNNs) (Li et al., 2021b, a; Zou et al., 2021), Graph Neural Networks (GNNs) (Chakraborty et al., 2021; Zhou et al., 2019; Nguyen et al., 2022; Steenhoek et al., 2024a), and Pre-trained Language Models (PLMs) (Hanif and Maffeis, 2022; Fu and Tantithamthavorn, 2022; Liu et al., 2024). These methods are relatively effective since they can learn vulnerability patterns in an end-to-end manner. Particularly, leveraging pre-trained knowledge, PLMs formulate vulnerability detection as a binary classification on code embedding, proven superior to RNNs and GNNs at detecting vulnerabilities (Steenhoek et al., 2023). Although PLMs show promise to some extent, these methods still face limitations in fully understanding code semantics for extracting representative or unseen vulnerability patterns (Ding et al., 2024b).
@@
Large Language Models (LLMs) have recently garnered widespread attention and shown strong potential across a variety of code-related tasks (Hou et al., 2023). Several empirical studies (Fu et al., 2023a; Zhou et al., 2024a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c) have explored using LLMs for vulnerability detection. For instance, Fu et al. (2023a) found that GPT-3.5-Turbo and GPT-4 fail to detect single-language (i.e., C/C++) vulnerabilities at both function-level and line-level.  Yin (2024) revealed that ChatGPT, when using prompts alone, can be easily swayed to change vulnerability classifications, indicating low confidence. Furthermore,  Yin et al. (2024) demonstrated that while fine-tuned LLMs can detect vulnerabilities, they perform more weakly than PLMs. Moreover, some LLM-based AVD techniques have been proposed, including Vul-RAG (Du et al., 2024), MSIVD (Yang et al., 2024a), and GRACE (Lu et al., 2024). These techniques demonstrate that LLMs are competitive with PLMs at detecting vulnerabilities in individual programming languages. For example, MSIVD applied multitask learning with LoRA to fine-tune the LLM and incorporated fine-tuned LLM embedding and Abstract Syntax Tree (AST) embedding. Vul-RAG checks code vulnerability by reasoning about vulnerability causes and fixing solutions from retrieved vulnerability knowledge, proving LLMs’ potential in detecting code vulnerabilities.
Despite significant advances in automated vulnerability detection using LLMs and PLMs, we identify three critical limitations in current approaches: (I) Existing research predominantly focuses on single-language vulnerability detection. Most existing studies limited their vulnerability detection capabilities to C and C++ using datasets like CVEfixes (Bhandari et al., 2021) and Big-Vul (Fan et al., 2020). This narrow focus fails to address the reality of modern development environments, which frequently incorporate multiple programming languages (Li et al., 2022). Studies have demonstrated the prevalence of vulnerabilities across diverse language ecosystems, including Python  (Alfadel et al., 2023) and Go (Hu et al., 2024). For AVD to be practically valuable, research must expand beyond single-language approaches to address multilingual vulnerability detection. In this context, we define multilingual detection not as analyzing mixed-language files simultaneously, but as the model’s ability to robustly generalize detection performance across independent languages with varying syntax and semantics. (II) The generalizability of PLMs and LLMs across diverse language-specific vulnerability profiles remains insufficiently explored. While PLMs and LLMs are pre-trained on vast multilingual corpora, it is unclear if this linguistic exposure translates to an understanding of vulnerability semantics across different paradigms. The nature of software vulnerabilities is heavily influenced by the underlying language paradigm. Low-level languages like C and C++ frequently exhibit memory corruption issues (e.g., buffer overflows) due to manual memory management, whereas high-level languages like Python tend to manifest vulnerabilities related to business logic or input injection. Existing study (Steenhoek et al., 2024b) often assume that pre-training equates to detection capability, failing to empirically validate whether models can effectively generalize vulnerability detection rules across these distinct language ecosystems, rather than simply memorizing syntax. This knowledge gap hinders effective model selection for systems that must secure software across a diverse technology stack. (III) Current evaluation frameworks lack a comprehensive assessment across different granularity levels. Vulnerability detection operates primarily at two levels: function-level and line-level. Function-level detection identifies whether a function contains vulnerabilities but lacks precision, potentially impeding developers from efficiently addressing specific issues. Conversely, line-level detection aims to precisely identify the vulnerable lines within a function. Existing research (Zhou et al., 2024a) has concentrated primarily on function-level detection in single-language contexts, leaving line-level multilingual vulnerability detection significantly underexplored. This gap is particularly problematic as multilingual contexts introduce greater challenges for line-level detection due to variations in syntax and semantics across programming languages, which complicate the effective generalization of language models.
@@
Our key findings are: (1) An appropriate LLM with instruction tuning and few-shot prompting demonstrates promising multilingual vulnerability detection performance. Specifically, for both function-level and line-level multilingual vulnerability detection, GPT-4o with instruction tuning and few-shot prompting demonstrates superior performance, achieving the highest accuracy (0.7196) at function-level compared to the best-performing CodeT5P (0.6037), and the best F1-score (0.6641) at line-level compared to other studied PLMs and LLMs. (2) Across different programming languages, the best-performing LLM achieves its highest function-level accuracy (0.8082) with Go and its lowest (0.6626) with Python. At the line-level, GPT-4o attains its highest F1-score (0.7815) on JavaScript, while its lowest F1-score (0.4348) is observed on C#. (3) Orthogonality analysis reveals that the best-performing LLM excels not only in uniquely correct detections but also shows higher accuracy in identifying the top-25 most dangerous CWE-IDs compared to PLMs. Base on CVSS severity, the top LLM also surpasses other studied PLMs and LLMs in detecting high-severity vulnerabilities (e.g., Critical and High level severity). (4) Furthermore, we investigate the impact of LLM size and compare effectiveness between reasoning and non-reasoning LLMs. The results show that the size of LLMs is not the decisive factor affecting performance on multilingual vulnerability detection, and the use of reasoning LLMs does not lead to significant improvements.
@@
  1. 1.
    This study systematically evaluates PLMs and LLMs for function- and line-level multilingual vulnerability detection across seven programming languages. It also compares various popular LLM strategies, including zero-shot prompting, retrieval-based few-shot prompting, and instruction tuning.
  2. 2.
@@
  4. 4.
    We further discuss the impact of reasoning capabilities and larger-scale LLMs on both function-level and line-level multilingual vulnerabilities, and analyze the deployment costs associated with LLMs and PLMs.
  5. 5.
@@
    We have made the replication package publicly available on our homepage (Shu, 2025), including all used data, codes, and analysis details involved in our study.
@@
Paper Extension. This paper extends our prior work on a preliminary study (Yu et al., 2025b), published as a research paper at the 1st International Workshop on Large Language Model Supply Chain Analysis, co-located with ISSTA’25. The key differences can be summarized as follows: (I) A broader range of studied models and prompting strategies. Specifically, we enhance PLMs by incorporating three OpenAI text-embedding models, and for LLMs, we explore two additional strategies: few-shot prompting and instruction tuning. (II) Incorporation of line-level detection. To provide a more comprehensive perspective, we extend the research scope beyond function-level detection to include line-level detection, enabling more detailed analysis of vulnerability locations. (III) Finer-grained analyses of detection performance. We conduct an in-depth analysis focusing on unique correct and incorrect detections, prediction tendencies for the Top-25 most dangerous CWE-IDs, and the effectiveness of detecting vulnerabilities across different severity levels. (IV) Exploration of effects of model size and architecture. We further investigate how the size of open-source LLMs and the use of reasoning-capable LLMs impact performance in multilingual vulnerability detection.
2 Study Design
@@ 2.1 Dataset Preparation
Studied dataset. To evaluate PLM-based AVD approaches and LLMs for multilingual vulnerability detection, we utilize the REEF dataset (Wang et al., 2023a). This dataset contains 4,466 CVEs with 30,987 patches across seven programming languages, including comprehensive vulnerability information (CVE, CWE, CVSS, etc.) and project details (such as commit messages). REEF is constructed from real-world vulnerabilities sourced from the National Vulnerability Database (NVD) and Mend’s CVE list (WhiteSource, 2022), an open-source vulnerability database covering 2016-2023. At the function-level, the dataset consists of 6,957 functions written in C, 2,244 in C++, 1,529 in C#, 3,187 in Go, 6,207 in Java, 5,066 in JavaScript, and 5,797 in Python. Our study encompasses all seven programming languages supported by REEF: C, C++, C#, Go, Java, JavaScript, and Python.
@@
Data pre-processing. To adapt the REEF dataset for multilingual AVD tasks, our first step was to extract the vulnerable and non-vulnerable functions. We received commit data in raw format with patches. To analyze both the original and modified versions of vulnerable functions, we processed these patches using the Linux patch command. [^1] https://www.man7.org/linux/man-pages/man1/patch.1.html Following the commit data collection (Fan et al., 2020), we pre-processed the data by removing code comments to minimize bias. Comments can contain misleading information, such as outdated or incorrect annotations. Additionally, they may include vulnerability-related details that pose a potential risk of data leakage. We then extracted function definition code using Tree-sitter (Brunsfeld, 2024), an efficient parser generator and incremental parsing library capable of analyzing multiple programming languages. To extract functions and their corresponding function-level and line-level labels, we iterated through function definitions in the pre-change file to match them with their post-change and pre-change function definitions. Specifically, the function names are used as identifiers to match and compare corresponding functions between the pre-change and post-change versions of the files. When multiple matching function definitions are found in the post-change file, we calculate the edit distance between the pre-change function and each post-change function. The pair with the smallest edit distance is then identified as the vulnerable and clean function pair. In other words, we define the pre-change function as the vulnerable function and the post-change function as the non-vulnerable function. Since many PLMs use the absolute position encoding, which is limited to the input length (e.g., 512), we filtered out functions that are greater than 512 in length. Finally, we obtained a total of 20,165 functions and corresponding labels. At the function-level, the dataset comprises 3,056 C, 1,792 C++, 427 C#, 2,905 Go, 3,235 Java, 5,468 JavaScript, and 3,282 Python functions. To obtain line-level labels, we first eliminated all non-vulnerable functions from the function-level dataset. We then applied DIFFLIB (Homepage, 2024) to match changed code lines by comparing each vulnerable function with its corresponding non-vulnerable version. DIFFLIB identifies three matching cases: (1) vulnerable functions fixed by adding code lines only, (2) vulnerable functions fixed by removing code lines only, and (3) vulnerable functions fixed by both removing and adding code lines. Since we need to locate specific line positions, only cases (2) and (3) can identify vulnerable line positions through deletion information. We use these two cases to obtain code line labels, i.e., line numbers and code lines, then filter out vulnerable functions where line labels cannot be extracted. To ensure the integrity of our line-level dataset, we excluded 9,613 functions from the training set, 1,202 from the validation set, and 1,216 from the test set because their vulnerable line positions could not be identified. Consequently, these samples were filtered out during the construction of our line-level dataset. Finally, we obtained a total of 8,134 vulnerable functions and the corresponding line labels. Finally, we obtained a total of 8,134 lines and corresponding labels. At the line-level, the dataset comprises 1,175 C, 651 C++, 151 C#, 919 Go, 1,375 Java, 2,496 JavaScript, and 1,367 Python functions.
To validate the robustness of Tree-sitter’s parsing capabilities, we conducted a manual validation on a statistically significant subset of 377 randomly selected samples (corresponding to a 95% confidence level with a 5% margin of error). To mitigate subjective bias, the first and third authors performed the verification independently. The first author validated 349 samples as correct (93% correctness rate), while the third author confirmed 346 (92% correctness rate). The resulting Cohen’s Kappa coefficient of 0.9448 indicates near-perfect agreement, providing strong evidence for Tree-sitter’s reliability. The validation followed a rigorous protocol: first, mapping code changes back to the original security patch hunks to isolate ground-truth vulnerable functions; second, cross-referencing these changes with pre- and post-patch function definitions to verify structural alignment and line-number consistency. Our inspection confirms that Tree-sitter provides a robust foundation for our parsing requirements.
@@
Table 1: Statistical summary of the function-level multilingual vulnerability dataset
@@
Table 2: Statistical summary of the line-level multilingual vulnerability dataset
@@
Construction of training and test datasets. Similar to Fu and Tantithamthavorn (2022), we divide the REEF dataset of function-level and line-level multilingual vulnerability into training, validation, and test sets at an 8:1:1 ratio. Specifically, we employed stratified random sampling to generate the data splits. Rather than performing a single global split, which risks under-representing minority languages, we partitioned the dataset into seven distinct strata, one for each language. Within each stratum, we applied the predetermined 8:1:1 ratio to generate language-specific training, validation, and test subsets. This stratification strategy serves two purposes: first, it ensures that every language is consistently represented across all pipeline stages; second, it preserves the natural variation in data volume across languages. By maintaining this skew, our dataset reflects the real-world prevalence of vulnerabilities in open-source repositories, ensuring that the model is evaluated on a distribution that aligns with the actual development landscape. As shown in Table 1, the function-level multilingual vulnerability dataset contains 16,126, 2,013, and 2,026 functions and corresponding labels for training, validation, and testing over seven languages. As shown in Table 2, the line-level multilingual vulnerability dataset contains 6,513, 811, and 810 vulnerable functions and corresponding line labels for training, validation, and testing across the studied languages.
@@
Figure 1: The severity distribution of multilingual vulnerability dataset
2.2 Experimented Language Models